A sample containing 5.60 g O 2 gas has a volume of 21.0 L . Pressure and temperature remain constant.Part A What is the new volume if 0.400 mole O 2 gas is added? Part B Oxygen is released until the volume is 7.50 L . How many moles of O 2 are removed? Part C What is the volume after 5.50 g He is added to the O 2 gas already in the container?
Answer #1

Part A : new volume = 69.0 L

Part B : moles O2 removed = 0.113 mol   (if wrong then try 0.513 mol)

Part C : volume = 186 L (if wrong then try 172 L)

Explanation

Part A

initial mass O2 = 5.60 g

initial moles O2 = (initial mass O2) / (moles O2)

initial moles O2 = (5.60 g) / (32.0 g/mol)

initial moles O2 = 0.175 mol

According to Avogadro's law,

V1 / n1 = V2 / n2

where V1 = initial volume = 21.0 L

n1 = initial moles O2 = 0.175 mol

V2 = final volume

n2 = final moles O2 = (0.175 mol) + (0.400 mol) = 0.575 mol

V2 = V1 * (n2 / n1)

V2 = (21.0 L) * (0.575 mol / 0.175 mol)

V2 = (21.0 L) * (3.28)

V2 = 69 L